OpenStudy (anonymous):

Write y = x + 5 in standard form using integers.

OpenStudy (anonymous):

Help Please

OpenStudy (anonymous):

@timo86m

OpenStudy (anonymous):

@RoseDryer

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

not at all i am trying to help my daughter with her homework and I don't know

OpenStudy (anonymous):

Standard form is in the form: Ax + By = C That basically means you are adding x and y together to get a constant value denoted "c." You must isolate v using basic algebra to do this. Subtract x from both sides of the equation to isolate the five. This will give you y - x = 5

OpenStudy (anonymous):

here are the choices A. –x – 6y = 30 B. 6x – y = 30 C. –x + 6y = 30 D. –x + 6y = 5

OpenStudy (anonymous):

None of those answer choices are possible with the given problem. Are you sure you listed the initial equation correctly? It asks for an integer answer, so that implies the original equation would have a non-integer coefficient somewhere. Is there a fraction before X or Y in the original equation ?

OpenStudy (skullpatrol):

Go to YouTube and search for Khan Academy equation of a line.

OpenStudy (skullpatrol):

http://www.youtube.com/watch?v=gvwKv6F69F0

OpenStudy (anonymous):

thanks